# Motives of nullcones of quiver representations

### v2

@@ 4.2. Rational/integral level quiver
For a quiver $Q$ we define the rational level quiver $Q_{\mathbb{Q}}$ with set of vertices $Q_{0}\times\mathbb{Q}$ and arrows
$(\alpha,a):(i,a)\rightarrow(j,a+1)$
for every arrow $\alpha:i\rightarrow j$ in $Q$ and every $a\in\mathbb{Q}$. This quiver obviously admits a decomposition into components $Q_{\mu}$ for $\mu\in[0,1[\cap\mathbb{Q}$, where $Q_{\mu}$ is the full subquiver of all vertices $(i,a)$ such that the fractional part $\{a\}$ of $a$ equals $\mu$.
We define a stability on $Q_{\mathbb{Q}}$ by $\theta(i,a)=-a$. We call a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ balanced if $\theta(\mathbf{e}|_{Q_{\mu}})=0$ for all $\mu\in[0,1[\cap\mathbb{Q}$ (where $\mathbf{e}|_{Q_{\mu}}$ denotes the restriction of $\mathbf{e}$ to the component $Q_{\mu}$). Every finitely supported dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ admits a projection $|\mathbf{e}|$ to a dimension vector for $Q$ by $|\mathbf{e}|_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}.$
@@
Using this notation, the index set ${\bar{L}_{\mathbf{d}}}$ for the Hesselink stratification of $N_{\mathbf{d}}(Q)$ is in bijection to the set of all balanced dimension vectors $\mathbf{e}$ for $Q_{\mathbb{Q}}$ such that $|\mathbf{e}|=\mathbf{d}$. Moreover, we have an isomorphism of semistable loci
$R^{\theta_{s}-sst}_{\mathbf{d}_{s}}(Q_{s})\cong R^{\theta-sst}_{\mathbf{e}}(Q_{\mathbb{Q}}).$
@@
For every entry $e_{i,a}$ in $\mathbf{e}$ we take $e_{i,a}$-times the entry $a$ in the coweight, in weakly ascending order for every vertex of $Q$. That this induces a bijection between balanced dimension vectors and balanced and dominant coweights now follows immediately from the definitions.
Now we have to compare the stabilities given by $\theta_{s}$ and $\theta$. Relabelling the vertices $(v,j)$ of the quiver $Q_{s}$ to $(v,p_{i}+j)$, we see directly that $Q_{s}$ can be viewed as a subquiver of the rational level quiver. Using the above construction of a balanced dominant coweight from a balanced dimension vector $\mathbf{e}$ it is then clear that, in both cases, the multiplicity of the entry $p_{i}+j$ (respectively $a$) in the string is encoded in the dimension vector. So the only essential difference between the stabilities is the factor $z$ of Definition 4.5. Since the set of semistable representations with respect to a stability $\theta$ does not change when passing from $\theta$ to $x\cdot\theta$ for positive $x$, it follows that, in fact, $R^{\theta_{s}-sst}_{\mathbf{d}_{s}}(Q_{s})\cong R^{\theta-sst}_{\mathbf{e}}(Q_{\mathbb{Q}})$. ∎
@@
The quiver $Q_{\mathbb{Q}}$ is the disjoint union of the $Q_{\mu}$ for $\mu\in[0,1[\cap\mathbb{Q}$, and every $Q_{\mu}$ can be identified with $Q_{\mu=0}=Q_{\mathbb{Z}}$, called the integral level quiver. We can reparametrize the index set $L^{\prime}_{\bf d}$: we map $\mathbf{e}$ to the tuple $(\mathbf{e}|_{Q_{\mu}})_{\mu}$ of dimension vectors, which we view as dimension vectors for $Q_{\mathbb{Z}}$. Now we interprete this reparametrization in terms of our stability conditions:
After identifying a balanced dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ with a tuple of dimension vectors $(\mathbf{e}_{\mu})_{\mu}$ for $Q_{\mathbb{Z}}$ we have:
$\theta(\mathbf{e}|_{Q_{\mu}})=\theta(\mathbf{e}_{\mu})-\mu\cdot\dim\mathbf{e}_{\mu}~{}.$
This condition can be rephrased using the slope function $\sigma=\theta/\dim$ and noting that $\theta(\mathbf{e}|_{\mathbb{Q}_{\mu}})=0$ since $\mathbf{e}$ is balanced:
$\sigma(\mathbf{e}_{\mu})=\frac{\theta(\mathbf{e}_{\mu})}{\dim\mathbf{e}_{\mu}}=\mu~{}$
@@
We can now state a final reformulation for the index set of Hesselink strata:
@@
5. Wall-crossing type formula
Our aim is to interprete the sum formula (2) as an ordered product formula for the corresponding generating series.
Describing the strata in terms of balanced dimension vectors $\mathbf{e}$ for the quiver $Q_{\mathbb{Q}}$, we can rewrite the formula
$\phi(s)=\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})-(\dim P_{s}-\dim G_{\mathbf{d}_{s}})$
in terms of $\mathbf{e}$.
Using Theorem 4.6, we see that $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$ is the dimension of the space of $|Q_{1}|$-tuples of $d\times d$-matrices with zero entries except possibly at position $(l,k)$ in the matrix corresponding to $\alpha:i\rightarrow j$ if $k\in I_{i}$, $l\in I_{j}$ and $s_{l}-s_{k}>1$. In terms of the rational level quiver this dimension is encoded in the dimensions $e_{i,a}$ at vertices which are not directly adjacent:
$\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})=\sum_{\alpha:i\rightarrow j}\sum_{a+1<b}e_{i,a}e_{j,b}~{}.$
We analyse the summand $\dim P_{s}-\dim G_{\mathbf{d}_{s}}$. The group $P_{s}$ was described in terms of matrix entries $(l,k)$ such that $k,l\in I_{i}$ for some vertex $i\in Q_{0}$ and $s_{l}-s_{k}\geq 0$, while $G_{\mathbf{d}_{s}}$ is determined by $s_{l}-s_{k}=0$. So we can construct $P_{s}$ from $G_{\mathbf{d}_{s}}$ by allowing arbitrary entries at positions $(l,k)$ where $k,l\in I_{i}$ for some vertex $i\in Q_{0}$ and $s_{l}-s_{k}>0$. For symmetry reasons the number of such positions is exactly
$\frac{\sum_{i\in Q_{0}}d_{i}^{2}-\dim G_{\mathbf{d}_{s}}}{2}~{}.$
@@ 5. Wall-crossing type formula
Alternatively we can note that $\dim P_{s}-\dim G_{\mathbf{d}_{s}}$ equals the dimension of the unipotent radical corresponding to the parabolic subgroup $P_{s}$.
With $d_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}$ and $\dim G_{\mathbf{d}_{s}}=\sum_{i\in Q_{0}}\sum_{a\in\mathbb{Q}}e_{i,a}^{2}$ we can conclude:
$\dim P_{s}-\dim G_{\mathbf{d}_{s}}=\frac{\sum_{i\in Q_{0}}(\sum_{a\in\mathbb{Q}}e_{i,a})^{2}-\sum_{i\in Q_{0}}\sum_{a\in\mathbb{Q}}e_{i,a}^{2}}{2}=\sum_{i\in Q_{0}}\sum_{a<b}e_{i,a}e_{i,b}$
yielding a formula for $\phi$ in terms of the dimension vector $\mathbf{e}$:
$\phi(\mathbf{e})=\sum_{\alpha:i\rightarrow j}\sum_{a+1<b}e_{i,a}e_{j,b}-\sum_{i\in Q_{0}}\sum_{a<b}e_{i,a}e_{i,b}~{}.$
@@
$\phi((\mathbf{e}_{\mu})_{\mu})$ $=\sum_{\alpha:i\rightarrow j}\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a-\mu}e^{\mu^{\prime}}_{j,b-\mu^{\prime}}-\sum_{i\in Q_{0}}\sum_{a<b}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a-\mu}e^{\mu^{\prime}}_{i,b-\mu^{\prime}}$
$=\sum_{\alpha:i\rightarrow j}\sum_{a+1+\mu<b+\mu^{\prime}}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a}e^{\mu^{\prime}}_{j,b}-\sum_{i\in Q_{0}}\sum_{a+\mu<b+\mu^{\prime}}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a}e^{\mu^{\prime}}_{i,b}$
The condition $a+1+\mu<b+\mu^{\prime}$ is equivalent to
$(\mu<\mu^{\prime}\text{ and }a+1=b)\text{ or }a+1<b$
for $\mu,\mu^{\prime}\in[0,1[$ and $a,b\in\mathbb{Z}$. Similarly, $a+\mu<b+\mu^{\prime}$ is equivalent to
$(\mu<\mu^{\prime}\text{ and }a=b)\text{ or }a<b.$
We thus obtain
$\phi((\mathbf{e}_{\mu})_{\mu})=$ $\sum_{\alpha:i\rightarrow j}(\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}+\sum_{a+1=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}})-$
$-\sum_{i}(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}})~{}.$
With our new parametrization of the index set and propositions 4.11 and 4.12, we can now state the following formula for the motive of $N_{\mathbf{d}}$:
@@
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}$ $=\sum_{s\in\tilde{L}_{\mathbf{d}}}\mathbb{L}^{\phi(s)}\frac{[R^{\theta_{s}-sst}_{\mathbf{d}_{s}}(Q_{s})]}{[G_{\mathbf{d}_{s}}]}$
$=\sum_{\mathbf{e}\in L^{\prime}_{\mathbf{d}}}\mathbb{L}^{\phi(\mathbf{e})}\frac{[R_{\mathbf{e}}^{\theta-sst}(Q_{\mathbb{Q}})]}{[G_{\mathbf{e}}]}$
$=\sum_{(\mathbf{e}_{\mu})_{\mu}\in L}\mathbb{L}^{\phi((\mathbf{e}_{\mu})_{\mu})}\prod_{\mu}\frac{[R_{\mathbf{e}_{\mu}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}~{}.$
@@
We define the $k$-shift ${\tau}^{k}\mathbf{e}$ of a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Z}}$ as the dimension vector
$({\tau}^{k}\mathbf{e})_{i,a}=\mathbf{e}_{i,a+k}~{}.$
@@
$-\langle\mathbf{e}^{\mu},{\tau}^{k}\mathbf{e}^{\mu^{\prime}}\rangle=-\sum_{i}\sum_{a}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{i\rightarrow j}\sum_{a}e_{i,a}^{\mu}e_{j,a+1+k}^{\mu^{\prime}}~{}.$
@@
$\sum_{\alpha:i\rightarrow j}(\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}+\sum_{a+1=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}})-$
$-\sum_{i}(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}).$
@@
$\sum_{\alpha:i\rightarrow j}\sum_{a\in\mathbb{Z}}(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,a+k+1}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,a+1}^{\mu^{\prime}})-$
$-\sum_{i}\sum_{a\in\mathbb{Z}}(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,a}^{\mu^{\prime}}),$
@@
$-\sum_{k>0}\sum_{\mu,\mu^{\prime}}\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle~{}.$
As before, we consider the ring of localized motives $R=K_{0}(\text{Var}_{\mathbb{C}})[\mathbb{L}^{-1},(1-\mathbb{L}^{i})^{-1},i\geq 1]$ and conclude from Theorem 5.2:
@@
In $R_{\mathbb{L}}\llbracket Q_{0}\rrbracket$, we can express the nilpotent motivic generating function
@@
$\sum\mathbb{L}^{-\sum_{k>0}\sum_{\mu,\mu^{\prime}\in[0,1[\cap\mathbb{Q}}\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle}\prod_{\mu\in[0,1[\cap\mathbb{Q}}\frac{[R_{\mathbf{e}_{\mu}}^{\theta-sst}(Q_{\mathbb{Z}})]}{[G_{\mathbf{e}_{\mu}}]}t^{|\sum_{\mu}\mathbf{e}_{\mu}|}~{},$
where the sum ranges over all sets $L({\bf d})$, that is, over all tuples $({\bf e}_{\mu})_{\mu\in[0,1[\cap\mathbb{Q}}$ of dimension vectors for $Q_{\mathbb{Z}}$ such that ${\bf e}_{\mu}=0$ for all but finitely many $\mu$, and $\sigma({\bf e}_{\mu})=\mu$ otherwise.
@@
We have seven tuples $({\bf e}_{\mu})_{\mu}$ such that $|\sum_{\mu}{\bf e}_{\mu}|={\bf d}$ for which semistable representations exist:
$({\bf(i,-1)}+{\bf(j,0)}+{\bf(k,1)})_{0},$
$({\bf(i,-1)}+{\bf(j,0)}+{\bf(k,0)})_{1/3},$
$({\bf(i,-1)}+{\bf(j,-1)}+{\bf(k,0)})_{2/3},$
$(({\bf(i,0)})_{0},({\bf(j,-1)}+{\bf(k,0)})_{1/2}),$
$(({\bf(j,0)})_{0},({\bf(i,-1)}+{\bf(k,0)})_{1/2}),$
$(({\bf(k,0)})_{0},({\bf(i,-1)}+{\bf(j,0)})_{1/2}),$
@@
$[N_{\bf d}]=(\mathbb{L}+2)(\mathbb{L}-1)^{2}+3(\mathbb{L}-1)+1=\mathbb{L}^{3},$

### v3

@@ 4.2 Rational/integral level quiver
For a quiver $Q$, we define the rational level quiver $Q_{\mathbb{Q}}$ with set of vertices $Q_{0}\times\mathbb{Q}$} and arrows $(\alpha,a)\colon(i,a)\rightarrow(j,a+1)$ for every arrow $\alpha\colon i\rightarrow j$} in $Q$ and every $a\in\mathbb{Q}$}. This quiver obviously admits a decomposition into components $Q_{\mu}$ for $\mu\in[0,1]\cap\mathbb{Q}$, where $Q_{\mu}$ is the full subquiver of all vertices $(i,a)$ such that the fractional part $\{a\}$ of $a$ equals $\mu$.
We define a stability on $Q_{\mathbb{Q}}$ by $\theta(i,a)=-a$. We call a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ balanced if $\theta(\mathbf{e}|_{Q_{\mu}})=0$ for all $\mu\in[0,1]\cap\mathbb{Q}$ (where $\mathbf{e}|_{Q_{\mu}}$ denotes the restriction of $\mathbf{e}$ to the component $Q_{\mu}$). Every finitely supported dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ admits a projection $|\mathbf{e}|$ to a dimension vector for $Q$ by $|\mathbf{e}|_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}.$
@@
Using this notation, the index set $\overline{L}_{\mathbf{d}}$} for the Hesselink stratification of $N_{\mathbf{d}}$ is in bijection to the set of all balanced dimension vectors $\mathbf{e}$ for $Q_{\mathbb{Q}}$ such that $|\mathbf{e}|=\mathbf{d}$. Moreover, we have an isomorphism of semistable loci $R^{\theta_{s}\text{\rm-sst}}_{\mathbf{d}_{s}}(Q_{s})\cong R^{\theta\text{\rm-sst}}_{\mathbf{e}}(Q_{\mathbb{Q}})$.
@@
For every entry $e_{i,a}$ in $\mathbf{e}$, we take $e_{i,a}$-times the entry $a$ in the coweight, in weakly ascending order for every vertex of $Q$. That this induces a bijection between balanced dimension vectors and balanced and dominant coweights now follows immediately from the definitions.
Now we have to compare the stabilities given by $\theta_{s}$ and $\theta$. Relabelling the vertices $(v,j)$ of the quiver $Q_{s}$ to $(v,p_{i}+j)$, we see directly that $Q_{s}$ can be viewed as a subquiver of the rational level quiver. Using the above construction of a balanced dominant coweight from a balanced dimension vector $\mathbf{e}$, it is then clear that in both cases the multiplicity of the entry $p_{i}+j$} (resp. $a$) in the string is encoded in the dimension vector. So the only essential difference between the stabilities is the factor $z$ of Definition 4.6. Since the set of semistable representations with respect to a stability $\theta$ does not change when passing from $\theta$ to $x\cdot\theta$ for positive $x$, it follows that, in fact, $R^{\theta_{s}\text{\rm-sst}}_{\mathbf{d}_{s}}(Q_{s})\cong R^{\theta\text{\rm-sst}}_{\mathbf{e}}(Q_{\mathbb{Q}})$. ∎
@@
The quiver $Q_{\mathbb{Q}}$ is the disjoint union of the $Q_{\mu}$ for $\mu\in[0,1]\cap\mathbb{Q}$, and every $Q_{\mu}$ can be identified with $Q_{\mu=0}=Q_{\mathbb{Z}}$, called the integral level quiver. We can reparametrize the index set $L^{\prime}_{\bf d}$: we map $\mathbf{e}$ to the tuple $(\mathbf{e}|_{Q_{\mu}})_{\mu}$ of dimension vectors, which we view as dimension vectors for $Q_{\mathbb{Z}}$. Now we interpret this reparametrization in terms of our stability conditions:
After identifying a balanced dimension vector $\mathbf{e}$ for $Q_{\mathbb{Q}}$ with a tuple of dimension vectors $(\mathbf{e}_{\mu})_{\mu}$ for $Q_{\mathbb{Z}}$, we have $\theta(\mathbf{e}|_{Q_{\mu}})=\theta(\mathbf{e}_{\mu})-\mu\cdot\dim\mathbf{e}_{\mu}$. This condition can be rephrased using the slope function $\sigma=\theta/{\dim}$ and noting that $\theta(\mathbf{e}|_{\mathbb{Q}_{\mu}})=0$ since $\mathbf{e}$ is balanced
$\sigma(\mathbf{e}_{\mu})=\frac{\theta(\mathbf{e}_{\mu})}{\dim\mathbf{e}_{\mu}}=\mu$
@@
We can now state a final reformulation for the index set of Hesselink strata.
@@
5 Wall-crossing type formula
Here we understand the notion of a wall-crossing formula in a purely formal sense: an identity of generating functions of motives involving an ordered product over slopes. A prototypical example is the identity [11, Theorem 2.4] expressing the generating series of motives of moduli stacks of quiver representations as an ordered product over such series for semistable quiver representations of fixed slope.
Our aim in this section is thus to interpret the sum formula (4.2) as an ordered product formula for the corresponding generating series.
Describing the strata in terms of balanced dimension vectors $\mathbf{e}$ for the quiver $Q_{\mathbb{Q}}$, we can rewrite the formula $\phi(s)=\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})-(\dim P_{s}-\dim G_{\mathbf{d}_{s}})$ in terms of $\mathbf{e}$.
Using Proposition 4.8, we see that $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$ is the dimension of the space of $|Q_{1}|$-tuples of $d\times d$ matrices with zero entries except possibly at position $(l,k)$ in the matrix corresponding to $\alpha\colon i\rightarrow j$ if $k\in I_{i}$, $l\in I_{j}$ and $s_{l}-s_{k}>1$. In terms of the rational level quiver, this dimension is encoded in the dimensions $e_{i,a}$ at vertices which are not directly adjacent
$\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})=\sum_{\alpha\colon i\rightarrow j}\sum_{a+1<b}e_{i,a}e_{j,b}.$
We analyse the summand $\dim P_{s}-\dim G_{\mathbf{d}_{s}}$. The group $P_{s}$ was described in terms of matrix entries $(l,k)$ such that $k,l\in I_{i}$ for some vertex $i\in Q_{0}$ and $s_{l}-s_{k}\geq 0$, while $G_{\mathbf{d}_{s}}$ is determined by $s_{l}-s_{k}=0$. So we can construct $P_{s}$ from $G_{\mathbf{d}_{s}}$ by allowing arbitrary entries at positions $(l,k)$ where $k,l\in I_{i}$ for some vertex $i\in Q_{0}$ and $s_{l}-s_{k}>0$. For symmetry reasons, the number of such positions is exactly
$\frac{\sum_{i\in Q_{0}}d_{i}^{2}-\dim G_{\mathbf{d}_{s}}}{2}.$
@@ 5 Wall-crossing type formula
Alternatively, we can note that $\dim P_{s}-\dim G_{\mathbf{d}_{s}}$ equals the dimension of the unipotent radical corresponding to the parabolic subgroup $P_{s}$.
With $d_{i}=\sum_{a\in\mathbb{Q}}e_{i,a}$ and $\dim G_{\mathbf{d}_{s}}=\sum_{i\in Q_{0}}\sum_{a\in\mathbb{Q}}e_{i,a}^{2}$, we can conclude
$\dim P_{s}-\dim G_{\mathbf{d}_{s}}=\frac{\sum_{i\in Q_{0}}\bigl(\sum_{a\in\mathbb{Q}}e_{i,a}\bigr)^{2}-\sum_{i\in Q_{0}}\sum_{a\in\mathbb{Q}}e_{i,a}^{2}}{2}=\sum_{i\in Q_{0}}\sum_{a<b}e_{i,a}e_{i,b}$
yielding a formula for $\phi$ in terms of the dimension vector $\mathbf{e}$
$\phi(\mathbf{e})=\sum_{\alpha\colon i\rightarrow j}\sum_{a+1<b}e_{i,a}e_{j,b}-\sum_{i\in Q_{0}}\sum_{a<b}e_{i,a}e_{i,b}.$
@@
$\phi((\mathbf{e}_{\mu})_{\mu})$ $=\sum_{\alpha\colon i\rightarrow j}\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a-\mu}e^{\mu^{\prime}}_{j,b-\mu^{\prime}}-\sum_{i\in Q_{0}}\sum_{a<b}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a-\mu}e^{\mu^{\prime}}_{i,b-\mu^{\prime}}$
$=\sum_{\alpha\colon i\rightarrow j}\sum_{a+1+\mu<b+\mu^{\prime}}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a}e^{\mu^{\prime}}_{j,b}-\sum_{i\in Q_{0}}\sum_{a+\mu<b+\mu^{\prime}}\sum_{\mu,\mu^{\prime}}e^{\mu}_{i,a}e^{\mu^{\prime}}_{i,b}.$
The condition $a+1+\mu<b+\mu^{\prime}$ is equivalent to ($\mu<\mu^{\prime}$ and $a+1=b$) or $a+1<b$ for $\mu,\mu^{\prime}\in[0,1]$} and $a,b\in\mathbb{Z}$. Similarly, $a+\mu<b+\mu^{\prime}$ is equivalent to ($\mu<\mu^{\prime}$ and $a=b$) or a<b{a<b}. We thus obtain
$\phi((\mathbf{e}_{\mu})_{\mu})={}$ $\sum_{\alpha\colon i\rightarrow j}\biggl(\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}+\sum_{a+1=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}\biggr)$
$-\sum_{i}\biggl(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}\biggr).$
With our new parametrization of the index set and Propositions 4.16 and 4.17, we can now state the following formula for the motive of $N_{\mathbf{d}}$.
@@
$\begin{split}\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}&=\sum_{s\in\overline{L}_{\mathbf{d}}}\mathbb{L}^{\phi(s)}\frac{\big[R^{\theta_{s}\text{\rm-sst}}_{\mathbf{d}_{s}}(Q_{s})\big]}{[G_{\mathbf{d}_{s}}]}=\sum_{\mathbf{e}\in L^{\prime}_{\mathbf{d}}}\mathbb{L}^{\phi(\mathbf{e})}\frac{\big[R_{\mathbf{e}}^{\theta\text{\rm-sst}}(Q_{\mathbb{Q}})\big]}{[G_{\mathbf{e}}]}\\ &=\sum_{(\mathbf{e}_{\mu})_{\mu}\in L(\mathbf{d})}\mathbb{L}^{\phi((\mathbf{e}_{\mu})_{\mu})}\prod_{\mu}\frac{\big[R_{\mathbf{e}_{\mu}}^{\theta\text{\rm-sst}}(Q_{\mathbb{Z}})\big]}{[G_{\mathbf{e}_{\mu}}]}.\end{split}$
@@
We define the $k$-shift ${\tau}^{k}\mathbf{e}$ of a dimension vector $\mathbf{e}$ for $Q_{\mathbb{Z}}$ as the dimension vector $\bigl({\tau}^{k}\mathbf{e}\bigr)_{i,a}=\mathbf{e}_{i,a+k}$.
@@
$-\big\langle\mathbf{e}^{\mu},{\tau}^{k}\mathbf{e}^{\mu^{\prime}}\big\rangle=-\sum_{i}\sum_{a}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{i\rightarrow j}\sum_{a}e_{i,a}^{\mu}e_{j,a+1+k}^{\mu^{\prime}}.$
@@
$\sum_{\alpha\colon i\rightarrow j}\biggl(\sum_{a+1<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}+\sum_{a+1=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,b}^{\mu^{\prime}}\biggr)$
$\qquad-\sum_{i}\biggl(\sum_{a<b}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}+\sum_{a=b}\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,b}^{\mu^{\prime}}\biggr).$
@@
$\sum_{\alpha\colon i\rightarrow j}\sum_{a\in\mathbb{Z}}\biggl(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{j,a+k+1}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{j,a+1}^{\mu^{\prime}}\biggr)$
$\qquad-\sum_{i}\sum_{a\in\mathbb{Z}}\biggl(\sum_{k>0}\sum_{\mu,\mu^{\prime}}e_{i,a}^{\mu}e_{i,a+k}^{\mu^{\prime}}+\sum_{\mu<\mu^{\prime}}e_{i,a}^{\mu}e_{i,a}^{\mu^{\prime}}\biggr),$
@@
$-\sum_{k>0}\sum_{\mu,\mu^{\prime}}\big\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\big\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle.$
As before, we consider the ring of localized motives $\mathbb{M}=K_{0}(\text{Var}_{\mathbb{C}})\big[\mathbb{L}^{-1},\bigl(1-\mathbb{L}^{i}\bigr)^{-1},i\geq 1\big]$ and conclude from Theorem 5.2 the following.
@@
In $\mathbb{M}\llbracket Q_{0}\rrbracket$, we can express the nilpotent motivic generating function
@@
$\sum\mathbb{L}^{-\sum_{k>0}\sum_{\mu,\mu^{\prime}\in[0,1]\cap\mathbb{Q}}\langle\mathbf{e}_{\mu},{\tau}^{k}\mathbf{e}_{\mu^{\prime}}\rangle-\sum_{\mu<\mu^{\prime}}\langle\mathbf{e}_{\mu},\mathbf{e}_{\mu^{\prime}}\rangle}\prod_{\mu\in[0,1]\cap\mathbb{Q}}\frac{\big[R_{\mathbf{e}_{\mu}}^{\theta\text{\rm-sst}}(Q_{\mathbb{Z}})\big]}{[G_{\mathbf{e}_{\mu}}]}t^{|\sum_{\mu}\mathbf{e}_{\mu}|},$
where the sum ranges over $L({\bf d})$, that is, over all tuples $({\bf e}_{\mu})_{\mu\in[0,1]\cap\mathbb{Q}}$ of dimension vectors for $Q_{\mathbb{Z}}$ such that ${\bf e}_{\mu}=0$ for all but finitely many $\mu$, and $\sigma({\bf e}_{\mu})=\mu$ otherwise.
@@
We have seven tuples $({\bf e}_{\mu})_{\mu}$ such that $|\sum_{\mu}{\bf e}_{\mu}|={\bf d}$ for which semistable representations exist
$({\bf(i,-1)}+{\bf(j,0)}+{\bf(k,1)})_{0},\qquad({\bf(i,-1)}+{\bf(j,0)}+{\bf(k,0)})_{1/3},$
$({\bf(i,-1)}+{\bf(j,-1)}+{\bf(k,0)})_{2/3},\qquad(({\bf(i,0)})_{0},({\bf(j,-1)}+{\bf(k,0)})_{1/2}),$
$(({\bf(j,0)})_{0},({\bf(i,-1)}+{\bf(k,0)})_{1/2}),\qquad(({\bf(k,0)})_{0},({\bf(i,-1)}+{\bf(j,0)})_{1/2}),$
@@
$N_{\bf d}]=(\mathbb{L}+2)(\mathbb{L}-1)^{2}+3(\mathbb{L}-1)+1=\mathbb{L}^{3},$